Stock XYZ has a current dividend of $7.00 . The dividend is expected to grow forever at a rate of 1.80% . Based on the riskiness of XYZ, its discount rate is 2.00% With this information, what is the dividend yield from today to period 1?

Group of answer choices 1.98% 0.22% 0.20% 1.80%
Answer #1

Answer: Correct option is 0.20%

Stock value=(Next year dividend)/(Discount rate-Growth rate)
=7*(1+1.8%)/(2%-1.8%)=3563
Dividend yield from today to year 1 is
=7/3563=0.001964637 or 0.20% (Rounded to two decimal places)